OpenStudy (anonymous):

What is true about the solutions of x2 − 2x − 15 = 0?

OpenStudy (anonymous):

No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

Using the quadratic formula. The discriminant \[b^2-4ac = (-2)^2-4(1)(-15)=64\] is greater than zero so there can't be no real solutions. So it must be either of the three remaining choices.

OpenStudy (anonymous):

Also the discriminant is not equal to zero so there can't be two identical rational solutions.

OpenStudy (anonymous):

You see a quadratic where \[b^2-4ac<0\] has no real solutions. A quadratic with \[b^2-4ac>0\] has 2 real or 2 irrational solutions. And a quadratic with \[b^2-4ac=0\] has two identical solutions.

OpenStudy (anonymous):

We are down to the two choices two different irrational solutions and two different rational solutions. To have rational solutions the discriminant must be a perfect square. And in my first post I calculated it to be 64 which is a perfect square. So the quadratic equation has two rational solutions.

OpenStudy (anonymous):

The answer is two rational solutions.

OpenStudy (anonymous):

I mean two different rational solutions